# Enhancing Suicide Risk Assessment: A Speech-Based Automated Approach in Emergency Medicine

### v1

@@ Abstract
The delayed access to specialized psychiatric assessments and care for patients at risk of suicidal tendencies in emergency departments creates a notable gap in timely intervention, hindering the provision of adequate mental health support during critical situations. To address this, we present a non-invasive, speech-based approach for automatic suicide risk assessment. For our study, we have collected a novel dataset of speech recordings from $20$ patients from which we extract three sets of features, including wav2vec, interpretable speech and acoustic features, and deep learning-based spectral representations. We proceed by conducting a binary classification to assess suicide risk in a leave-one-subject-out fashion. Our most effective speech model achieves a balanced accuracy of $66.2\,\%$. Moreover, we show that integrating our speech model with a series of patients’ metadata, such as the history of suicide attempts or access to firearms, improves the overall result. The metadata integration yields a balanced accuracy of $94.4\,\%$, marking an absolute improvement of $28.2\,\%$, demonstrating the efficacy of our proposed approaches for automatic suicide risk assessment in emergency medicine.
@@ I INTRODUCTION
However, only some works analyzed speech to predict suicidality. Building upon the observation that the characteristics of the patient’s voice are more important than the content of what was spoken, Ozdas et al. [ozdas2004analysis] used Mel-cepstral coefficients in conjunction with Gaussian mixture models to predict near-term suicidal risk. Cummins et al. [cummins2015review] further analyzed how common paralinguistic speech characteristics are affected by suicidality.
@@ II DATASET
We have collected a unique speech-based dataset from $20$ subjects undergoing emergency admission to the Department of Psychiatry at the University of Augsburg, Germany. These recordings encompass three distinct types of speech activities: (i) reading of three short narratives [^2] The three short stories were (1) “Der Nordwind und Die Sonne” (the story of ‘the north wind and the sun’), “Gleich am Walde” (the story of ‘right next to the forest’), and “Der Hund und das Stück Fleisch” (the story of ‘the dog and the piece of meat’). widely used within phonetics and speech pathology, each story was read twice; (ii) picture description where each patient describes a picture in their words; and (iii) isolated vowels production (/a:/, /e:/, /i:/, /o:/, /u:/). Each subject contributed 8 recordings after admission to the emergency room (referred to as baseline), and some consented to record an additional 8 upon discharge from hospital care. In our experiments, we restrict the analysis to the baseline samples to allow for a consistent validation procedure. The study doctor assessed the suicide risk level of each patient using the Likert scale [1 – 6].
@@ III-B1 Speech and Acoustic Functionals
We utilise openSMILE [Eyben10-OTM] to extract two sets of features – a comprehensive set of 6 37363736\,3736 373 functionals (ComParE) and eGeMAPS, a smaller selection of features (88 dimensional) crafted for speech emotion recognition.
@@ III-D Experimental Setup
We conduct binary classification of suicide risk based on assessments from the study doctor. Both physicians rate suicidality on a scale ranging from 1 (not suicidal) to 6 (acute suicide attempt). These ratings are then categorized into classes representing low suicidality (1-4) and high risk of suicide (5 and 6), forming the targets for our classification models. We evaluate these models using a leave-one-subject-out (LOSO) validation scheme. Classification is performed by fitting linear Support Vector Machine (SVM) models on the extracted audio features and vectorized metadata, with categorical features like gender encoded using one-hot encoding. The data is normalized to zero mean and unit standard deviation based on each fold’s training partition. We optimize the SVM’s regularisation constant on a logarithmic scale from $1$ to $10^{-7}$ via an inner 5-fold cross-validation. The evaluation metric is balanced accuracy, which accounts for class imbalances by considering the average accuracy of each class.
@@ V CONCLUSIONS AND FUTURE WORK
Future work includes expanding the dataset by incorporating a larger cohort of individuals with diverse demographic backgrounds. Additionally, novel metadata sources, such as social media activity, or lifestyle habits could offer complementary information that enhances the accuracy of our speech models. Finally, as proposed in [gerczuk2023zero], integrating longitudinal data to track changes in speech patterns and risk factors over time could enable the development of dynamic risk assessment models capable of monitoring and predicting evolving suicidal tendencies.

### v2

# Non-Invasive Suicide Risk Prediction Through Speech Analysis

@@ Abstract
The delayed access to specialized psychiatric assessments and care for patients at risk of suicidal tendencies in emergency departments creates a notable gap in timely intervention, hindering the provision of adequate mental health support during critical situations. To address this, we present a non-invasive, speech-based approach for automatic suicide risk assessment. For our study, we collected a novel speech recording dataset from $20$ patients. We extract three sets of features, including wav2vec, interpretable speech and acoustic features, and deep learning-based spectral representations. We proceed by conducting a binary classification to assess suicide risk in a leave-one-subject-out fashion. Our most effective speech model achieves a balanced accuracy of $66.2\,\%$. Moreover, we show that integrating our speech model with a series of patients’ metadata, such as the history of suicide attempts or access to firearms, improves the overall result. The metadata integration yields a balanced accuracy of $94.4\,\%$, marking an absolute improvement of $28.2\,\%$, demonstrating the efficacy of our proposed approaches for automatic suicide risk assessment in emergency medicine.
@@ I INTRODUCTION
[…]
@@ II DATASET
We have collected a unique speech-based dataset from $20$ subjects undergoing emergency admission to the Department of Psychiatry at the BLINDED FOR REVIEW. These recordings encompass three distinct types of speech activities: (i) reading of three short narratives [^2] The three short stories were (1) “Der Nordwind und Die Sonne” (the story of ‘the north wind and the sun’), “Gleich am Walde” (the story of ‘right next to the forest’), and “Der Hund und das Stück Fleisch” (the story of ‘the dog and the piece of meat’). widely used within phonetics and speech pathology, each story was read twice; (ii) picture description where each patient describes a picture in their words; and (iii) isolated vowels production (/a:/, /e:/, /i:/, /o:/, /u:/). Each subject contributed 8 recordings after admission to the emergency room (referred to as baseline), and some consented to record an additional 8 upon discharge from hospital care. In our experiments, we restrict the analysis to the baseline samples to allow for a consistent validation procedure. The study doctor assessed the suicide risk level of each patient using the Likert scale [1 – 6].
@@ III-B1 Speech and Acoustic Functionals
We utilize openSMILE [Eyben10-OTM] to extract two sets of features – a comprehensive set of 6 37363736\,3736 373 functionals (ComParE) and eGeMAPS, a smaller selection of features (88 dimensional) crafted for speech emotion recognition.
@@ III-D Experimental Setup
We conduct binary classification of suicide risk based on assessments from the study doctor. Both physicians rate suicidality on a scale ranging from 1 (not suicidal) to 6 (acute suicide attempt). These ratings are then categorized into classes representing low suicidality (1-4) and high risk of suicide (5 and 6), forming the targets for our classification models. We evaluate these models using a leave-one-subject-out (LOSO) validation scheme. Classification is performed by fitting linear Support Vector Machine (SVM) models on the extracted audio features and vectorized metadata, with categorical features like gender encoded using one-hot encoding. The data is normalized to zero mean and unit standard deviation based on each fold’s training partition. We optimize the SVM’s regularization constant on a logarithmic scale from $1$ to $10^{-7}$ via an inner 5-fold cross-validation. The evaluation metric is balanced accuracy, which accounts for class imbalances by considering the average accuracy of each class.
@@ V CONCLUSIONS AND FUTURE WORK
Future work involves expanding the dataset to a larger, more diverse cohort and improving model generalization. Additional metadata like social media activity or lifestyle habits could enhance accuracy. Incorporating longitudinal data to track speech pattern changes may enable dynamic risk assessment of suicidal tendencies [gerczuk2023zero]. Finally, using advanced affective computing models like ExHuBERT [^3] https://huggingface.co/amiriparian/ExHuBERT [Amiriparian24-EEH] could offer deeper emotional analysis, further refining risk prediction.